# VAMP-Net: An Interpretable Multi-Path Framework of Genomic Permutation-Invariant Set Attention and Quality-Aware 1D-CNN for MTB Drug Resistance

### v1

@@ 6 Conclusion
We have introduced the Interpretable Variant-Aware Multi-Path Network (VAMP-Net), a novel deep learning framework that successfully addresses two critical, often-overlooked challenges in genomic machine learning: modeling genetic epistasis and systematically incorporating variant calling quality into the prediction process. Our work delivers a significant methodological advance that provides both state-of-the-art predictive performance and crucial layers of auditable interpretability, essential for bridging computational genomics with clinical utility.
The core of VAMP-Net lies in its dual-path architecture and the Adaptive Fusion mechanism. The Set Attention Transformer (Path-1) proved highly effective by operating on the inherently permutation-invariant set of ChromPos_Ref>>Alt variant tokens, successfully elucidating complex genetic epistasis and discovering both canonical resistance drivers and influential non-canonical loci. The Integrated Gradients attribution analysis provides a clear, ranked list of mutations and their co-dependencies, forming a direct basis for targeted functional validation.
However, the most compelling finding lies in the contribution of the Quality-Aware pathway. VAMP-Net is the first model to systematically integrate variant calling quality into resistance prediction, addressing the critical gap where existing methods treat all variant calls equally regardless of technical confidence. Our quantitative feature importance scores demonstrate that the fusion architecture performs a learned feature calibration and noise regularization . Specifically, the pronounced increase in reliance on VCF quality metrics (e.g., FRS, GT_CONF_PERCENTILE) for the more complex Rifabutin (RFB) phenotype, compared to Rifampicin (RIF), confirms the fusion mechanism intelligently and dynamically adapts its weighting based on the complexity and confidence requirements of the task. This dynamic weighting mechanism provides an unprecedented level of robustness.
The VAMP-Net framework’s fundamental design is inherently generalizable and extends well beyond the context of Multi-drug resistance in M. tuberculosis. Our proposal can be applied to any genotype-phenotype problem characterized by three conditions:
  • Variable-Length Inputs: The sample is defined by an unordered set of symbolic events (e.g., structural variants, somatic mutations in cancer, or transcriptomic changes).
  • Epistatic Dependencies: The final phenotype is driven by complex interactions rather than single-gene effects.
  • Heterogeneous Evidence: The symbolic input (genotype) is accompanied by quantitative, context-specific technical metadata (e.g., sequencing depth, coverage, or population frequency).
Future work will thus focus on extending this multi-path fusion paradigm to incorporate additional modalities in other clinical domains, such as incorporating image modalities (e.g., microscopy or histology) or longitudinal clinical metadata alongside genomic data. This work establishes a new standard for confidence and transparency in clinical AI systems by ensuring that all predictions are not only accurate but are driven by a mechanism that is transparent and auditable across both biological and technical evidence streams.

### v2

@@ 5 Conclusion
This study introduced the Interpretable Variant-Aware Multi-Path Network (VAMP-Net), a model designed to bridge the gap between high-dimensional genomic pathogenicity and technical data reliability. By integrating a Permuutation-Invariant Set Attention Transformer for epistatic modeling with a Quality-Aware 1D-CNN, VAMP-Net addresses the critical challenge of separating biological signals from sequencing noise. Our results demonstrate that the dual-path architecture successfully recovers canonical resistance markers while discovering novel loci, such as mmuM and mmpL5, within statistically significant metabolic modules centered on cell-wall remodeling. The most compelling methodological advancement is the model’s capacity for adaptive feature calibration. Quantitative attribution analysis confirms that the fusion mechanism dynamically modulates its reliance on quality metrics (e.g., FRS and GT_CONF_PERCENTILE) based on the phenotypic complexity, providing an unprecedented level of decision robustness for subtle resistance profiles like Rifabutin. This "integrated audit" ensures that all predictions are supported by both biological and technical evidence streams, establishing a new standard for transparency in clinical AI.
Future work will focus on extending this multi-path paradigm to multi-modal clinical settings. By incorporating longitudinal metadata and medical imaging (e.g., histology or microscopy) alongside genomic sets, we aim to evolve VAMP-Net into a comprehensive diagnostic engine. Ultimately, this work provides a scalable, mathematically grounded template for genotype-phenotype association in any domain characterized by variable-length inputs, complex epistasis, and heterogeneous data quality.
Acknowledgement
This work was carried out partially using the AI Datacenter of the National School of Artificial Intelligence, funded under grant number E049 24 0117 by the Algerian Ministry of Higher Education and Scientific Research.
@@ Acknowledgement
Code availability
The complete code of this study can be found at https://github.com/Aicha-AI/VAMP-Net.MTB-Drug-Resistance/tree/main.